A truck is moving on a frictionless surface with uniform velocity of 10m/s.A leak occurs in the water tank of the truck at the rate of 2kg/s.What is the speed of truck after 50s if the mass of truck is 100kg and mass of water initially was 100kg?

Grade:

1 Answers

sanki kurli
13 years ago

initially the mass of the system =mass of ( truck+water)=100+100= 200kg

water is leaking at 2 kg/s

so, after 50s .... 2*50kg of water has leaked out.

so final mass of system= 200 - 100 =100kg

now conservation of momentum...

200* v= 100*x   ...x= final speed

=>x=2v.